OpenStudy (prettylady1210):

Graph the system of inequalities. Show by shading in the feasible region 4x+y <2 , y> - 2

jhonyy9 (jhonyy9):

like a first step 4x+y<2 => y < -4x+2 y>-2 do you can continue it ?

OpenStudy (sshayer):

first draw the line 4x+y=2 when x=0,y=2 when y=-2,4x=2+2=4,x=1 so draw a line through (0,2) and (1,-2) put x=0,y=0 in 4x+y<2 0+0<2, 0<2 which is true. Hence (0,0) lies on the graph. y>-2 put (0,0) in it. 0>-2 hence (0,0) lies on the graph. draw the line y=-2